RECENT PROGRESS IN THE NEAR INFRARED SPECTROSCOPY OF CO_{2}^{a}

\underline{CHARLES E. MILLER}, MATTHEW A. MONTGOMERY, ROBERT M. ONORATO, AND CHRISTOPHER R. JOHNSTONE, Department of Chemistry, Haverford College, Haverford, PA 19041-1392 USA; LINDA R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109-8099 USA.

We have begun a systematic reinvestigation of the CO_{2} spectrum in the 4000 - 9000 cm^{−1} range using the McMath-Pierce Fourier transform spectrometer at the Kitt Peak National Solar Observatory. Fits to more than 100 vibrational bands illustrate that it is possible to obtain rms uncertainties as low as 2 × 10^{−5} cm^{−1} for strong bands and uncertainties better than 2 × 10^{−4} for all bands with S_{v}^{0} > 0.01×10^{−22} cm^{−1}/(molecule cm^{−2}). We will also present the first analysis of several ^{18}O^{12}C^{18}O and ^{16}O^{13}C^{18}O bands. All results will be compared to the data available in the HITRAN database and evaluated relative to the needs of atmospheric remote sensing.

aPart of the research described in this paper was carried out by the Jet Propulsion Laboratory,California Institute of Technology, under contract with the National Aeronautics and Space Administration.